I have x $5 notes and y $10 notes. There are eight notes altogether and their total value is $55. How many of each note do i have?

Answer by checkley79(3341)

X+Y=8
X=8-Y
5X+10Y=55
5(8-Y)+10Y=55
40-5Y+10Y=55
5Y=55-40
5Y=15
Y=15/5Y
Y=3 10 DOLLAR BILLS.
X+3=8
X=8-3
X=5 FIVE DOLLAR BILLS.
PROOF:
5*5+3*10=55
25+30=55
55=55